A company can use cotton to produce gloves, hats, and shirts. The profit on a pair of gloves(G) is $5, the profit on a hat(H) is $9, and the profit on a shirt(S) is $10. The items are produced according to the following production functions:G = 36CGH = 23CH - 0.5CH2S = 80CS - CS2where CG, CH, and CS are the pounds of cotton.There are 120 pounds of cotton available. How many pounds of cotton should be used to make gloves (CG) ?Enter as a value.
Answer #1

Answer 1 Maximize : t = Total Profit = 59 +9H+ IOS Subject to : CH + CG + Cs = 120 (1) Now Ipe54 +911 7 105 - Thy = 5(364) +